# Sequential Monte-Carlo testing by betting

### v4

@@ Appendix B Calibrating $\mathfrak{p}^{\mathrm{perm}}_{B}$ into an e-value
Analogously, we can define p-to-e calibrators (in the following, we just call these calibrators). However, in this case there are many more possible choices of calibrators. Below, we show that classical calibrators are inadmissible in our setting and introduce a new class of admissible calibrators for such problems. Furthermore, since calibrators define betting strategies, we show that classical calibrators lead to strategies with low power by means of simulations.
@@
$\mathbb{E}[h(\mathfrak{p}^{\mathrm{perm}}_{B})]=\frac{1}{3}\frac{1}{2}\sqrt{3}+\frac{1}{3}\frac{1}{2}\sqrt{\frac{3}{2}}+\frac{1}{3}\frac{1}{2}\sqrt{1}=\frac{1}{6}\left(\sqrt{3}+\sqrt{\frac{3}{2}}+1\right)<1,$
which shows that $h$ is inadmissible in this case. In the following we provide a characterization of admissible calibrators for p-values with the same support as $\mathfrak{p}^{\mathrm{perm}}_{B}$.
@@
Let $\mathfrak{p}^{\mathrm{perm}}_{B}$ be a p-value with support $\{1/(B+1),\ldots,B/(B+1),1\}$ and $\mathbb{P}(\mathfrak{p}^{\mathrm{perm}}_{B}\leq r/(B+1))=r/(B+1)$ for all $11\leq r\leq B+$ 1. The calibrator $h$ is admissible for $\mathfrak{p}^{\mathrm{perm}}_{B}$ if and only if $\sum_{r=1}^{B+1}h(r/(B+1))=B+1$.
@@
The proposition follows immediately from $\mathbb{E}[h(\mathfrak{p}^{\mathrm{perm}}_{B})]=\sum_{r=1}^{B+1}h(r/(B+1))/(B+1)$ under the null hypothesis. ∎
An example of an admissible calibrator for $\mathfrak{p}^{\mathrm{perm}}_{B}$ is
$h(\mathfrak{p}^{\mathrm{perm}}_{B})=h(r/(B+1))=(B+1)/(r\ell_{B+1}),\quad 1\leq r\leq B+1,$ (16)
where $\ell_{B}=1+1/2+1/3+\dots 1/B$ is the $B$-th harmonic number. This is reminiscent of the Benjamini-Yekutieli correction in multiple testing [1]. Indeed, if these e-values (calibrated p-values) are plugged into the e-BH procedure [33], one obtains a method that is equivalent to applying the BY procedure on the original p-values.
@@
$h(\mathfrak{p}^{\mathrm{perm}}_{B})=h(r/(B+1))=\frac{(B+1)/\sqrt{r}}{s_{B+1}}>\sqrt{B+1}/(2\sqrt{r})=1/(2\sqrt{\mathfrak{p}^{\mathrm{perm}}_{B}}),$ (17)
@@
where $s_{B+1}=\sum_{i=1}^{B+1}1/\sqrt{i}<2\sqrt{B+1}-1$, which shows that this is a uniform improvement of the calibrator by Vovk and Wang [31] for $\kappa=1/2$.
Note that $R_{B}/(B+1)$, where $R_{B}$ is the rank of the $B$-th permutation, is also a valid p-value under the null hypothesis with the same support as $\mathfrak{p}^{\mathrm{perm}}_{B}$, even conditional on the ranks of the previous permutations $R_{1}^{B-1}$. Also note that the condition derived in Proposition B.1 is exactly the same as we derived for our betting strategies in Section 1.1. Therefore, every admissible calibrator $h$ for $\mathfrak{p}^{\mathrm{perm}}_{B}$ defines a bet $B_{B}$ on the rank $R_{B}$ and vice versa. Hence, one approach to find betting strategies would be to just use the calibrators from 16 or 17. Note that the calibrators should be applied to the “p-value” $(B-R_{B}+2)/(B+1)$ instead of $R_{B}/(B+1)$, such that small values are evidence for the alternative.
In Figure 7, the power and number of permutations is displayed for the betting strategies obtained by the calibrators from  16 and 17 and for our binomial strategy. We used the same simulation setup as in Section 7. It is easy to see that the power of the calibration based strategies is very low. This exemplifies the importance of adapting the betting strategy to the previous ranks.
Figure 7: Power and number of permutations until the decision was obtained for $\alpha=0.05$ and different strengths of the alternative $\mu$.
Appendix C Using a Beta distribution as prior for $\mathfrak{p}^{\mathrm{lim}}$
@@
The function $\mathrm{Beta}(a,b)$ is called beta function. In this section we derive the wealth obtained by applying the binomial mixture strategy with a beta distribution as prior for $\mathfrak{p}^{\mathrm{lim}}$ and show some simulation results for this strategy.
Proposition C.1.
After $B$ permutations and $k$ losses, the wealth obtained by the binomial mixture strategy with density $g_{a,b}$ is given by $W_{B}^{g_{a,b}}(k)=\mathrm{Beta}(a+k,b+B-k)/(\mathrm{Beta}(k+1,B-k+1)\mathrm{Beta}(a,b))$.
@@
First note that $(B+1){B\choose k}=\mathrm{Beta}(k+1,B+1)$. Therefore, the wealth is given by
@@
$=\frac{\mathrm{Beta}(k+a,B+b-k)}{\mathrm{Beta}(k+1,B-k+1)\mathrm{Beta}(a,b)}.$
It can be shown that if $a\leq 1$ and $b\geq 1$, then $g_{a,b}(x)$ is nonincreasing in $x$ which seems reasonable as prior for $\mathfrak{p}^{\mathrm{lim}}$. In particular, for $a=1$ and $b\in\mathbb{N}$, the wealth simplifies to
$W_{B}^{g_{1,b}}(k)=b\frac{(B+1)\cdots(B-k+1)}{(B+b)\cdots(B-k+b)},$
which is decreasing for an increasing number of losses. In Figure 8 the log-transformed p-values obtained by the binomial mixture strategy with density $g_{1,b}$ are compared with the ones obtained by the binomial strategy, the binomial mixture strategy (with density $u_{0.9\alpha}$), the aggressive strategy and the classical permutation p-value. We used the same simulation setup as described in Section 7. The binomial mixture strategy with beta prior bets more aggressively when $b$ is large. However, the binomial strategy and the binomial mixture strategy with uniform prior lead to more rejections, as seen by the larger number of log-transformed p-values below the grey line $\log(\alpha)$, in all cases.
Figure 8: All plots display log-transformed p-values in ascending order for $2000$ simulations — following the experimental protocol in Subsection C. The grey horizontal line equals $\log(\alpha)$, where $\alpha$ equals 0.05. The binomial mixture strategy with a beta prior bets more aggressive when $b$ is large. However, in terms of rejections (p-values below grey line) it is outperformed by the binomial strategy and the binomial mixture strategy with uniform prior in all cases.

### v5

@@ Appendix B Calibrating $Q^{\mathrm{perm}}_{T}$ into an e-value
Analogously, we can define p-to-e calibrators (in the following, we just call these calibrators). However, in this case there are many more possible choices of calibrators. Below, we show that classical calibrators are inadmissible in our setting and introduce a new class of admissible calibrators for such problems.
@@
$\mathbb{E}[h(Q^{\mathrm{perm}}_{T})]=\frac{1}{3}\frac{1}{2}\sqrt{3}+\frac{1}{3}\frac{1}{2}\sqrt{\frac{3}{2}}+\frac{1}{3}\frac{1}{2}\sqrt{1}=\frac{1}{6}\left(\sqrt{3}+\sqrt{\frac{3}{2}}+1\right)<1,$
which shows that $h$ is inadmissible in this case. In the following we provide a characterization of admissible calibrators for p-values with the same support as $Q^{\mathrm{perm}}_{T}$.
@@
Let $Q^{\mathrm{perm}}_{T}$ be a p-value with support $\{1/(T+1),\ldots,T/(T+1),1\}$ and $\mathbb{P}(Q^{\mathrm{perm}}_{T}\leq r/(T+1))=r/(T+1)$ for all $11\leq r\leq T+$ 1. The calibrator $h$ is admissible for $Q^{\mathrm{perm}}_{T}$ if and only if $\sum_{r=1}^{T+1}h(r/(T+1))=T+1$.
@@
The proposition follows immediately from $\mathbb{E}[h(Q^{\mathrm{perm}}_{T})]=\sum_{r=1}^{T+1}h(r/(T+1))/(T+1)$ under the null hypothesis. ∎
An example of an admissible calibrator for $Q^{\mathrm{perm}}_{T}$ is
$h(Q^{\mathrm{perm}}_{T})=h(r/(T+1))=(T+1)/(rv_{T+1}),\quad 1\leq r\leq T+1,$ (20)
where $v_{T}=1+1/2+1/3+\dots 1/T$ is the $T$-th harmonic number. This is reminiscent of the Benjamini-Yekutieli correction in multiple testing [1]. Indeed, if these e-values (calibrated p-values) are plugged into the e-BH procedure [40], one obtains a method that is equivalent to applying the BY procedure on the original p-values.
@@
$h(Q^{\mathrm{perm}}_{T})=h(r/(T+1))=\frac{(T+1)/\sqrt{r}}{s_{T+1}}>\sqrt{T+1}/(2\sqrt{r})=1/(2\sqrt{Q^{\mathrm{perm}}_{T}}),$ (21)
@@
where $s_{T+1}=\sum_{i=1}^{T+1}1/\sqrt{i}<2\sqrt{T+1}-1$, which shows that this is a uniform improvement of the calibrator by Vovk and Wang [37] for $\kappa=1/2$.
Appendix C The Besag-Clifford p-value for $T=\infty$: a negative binomial permutation test
The Besag-Clifford strategy also offers an exact p-value if we do not specify an upper bound $T$ of maximum permutations, meaning that the p-value
$Q_{\gamma(h)},\quad\text{where }\gamma(h):=\lim_{T\to\infty}\gamma(h,T),$
is also valid and exact. While Besag and Clifford pointed that out, they, and also most follow up work, were mainly concerned with the case $T<\infty$ in order to obtain a closed (bounded) sampling scheme. Nevertheless, the maximum number of permutations drawn is still almost surely finite, and, conditional on the data, has finite expectation and variance. Indeed, conditional on $Q_{\mathrm{lim}}=q_{\mathrm{lim}}$, $\gamma(r)$ is just a sum of (shifted) geometric random variables, or equivalently, $\gamma(r)-r$ is negative binomial with “success” probability $q_{\mathrm{lim}}$, so we call the test as the negative binomial permutation test and provide the result below for easier reference.
Proposition C.1 (Negative binomial permutation test).
For any pre-defined number of losses $h\geq 1$, define $\gamma(h):=\inf\{t\geq 1:\sum_{i=1}^{t}\mathbbm{1}\{Y_{t}\geq Y_{0}\}\geq h\}$ as being the time of the $h$-th loss. Then, $Q_{\gamma(h)}:=h/\gamma(h)$ is a valid and exact p-value for $H_{0}$: for any $N\geq 0$, $\mathbb{P}_{H_{0}}(Q_{\gamma(h)}\leq h/(h+N))=h/(h+N)$.
Proof.
Note that $\mathrm{supp}(Q_{\gamma(h)})=\{1,h/(h+1),h/(h+2),\ldots\}$ and for every $N\in\mathbb{N}_{0}$, we have $\mathbb{P}_{H_{0}}(Q_{\gamma(h)}\leq h/(h+N))=\mathbb{P}_{H_{0}}(\gamma(h)>h+N-1)=h/(h+N)$, where the last equation follows from the fact that $\gamma(h)>h+N-1$ means that $Y_{0}$ is among the $h$ largest test statistics of $Y_{0},Y_{1},\ldots,Y_{h+N-1}$. ∎
We remark that $Q_{\gamma(h)}$ bears an interesting resemblance to the standard permutation p-value $Q^{\mathrm{perm}}_{T}$. Both of them equal the number of losses divided by the number of permutations; $Q^{\mathrm{perm}}_{T}$ fixes the denominator to equal $T$ and lets the numerator be determined by the data, while $Q_{\gamma(h)}$ fixes the numerator as $h$, and lets the denominator be determined by the data. The fact most relevant for us is that since
$Q^{\mathrm{agg}}_{\tau}=Q^{\mathrm{agg}}_{\gamma(1)},$
the same exactness of p-value holds for $Q^{\mathrm{agg}}_{\tau}$. Said differently, one can obtain a valid and exact p-value as the inverse of the first time at which a permuted statistic exceeds the original unpermuted statistic. We summarize this in another proposition for ease of future reference.
Proposition C.2.
If the aggressive strategy is stopped when its wealth hits zero, as defined by the time $\gamma(1)$, the resulting stopped p-value equals $1/\gamma(1)$, and is an exact p-value. Further, if one stops at some arbitrary stopping time $\rho\leq\gamma(1)$, then $1/\rho$ is also a valid (potentially inexact) p-value.
Thus, if a level $\alpha$ is prespecified, one can stop as soon as $1/\rho$ drops below $\alpha$ (or the first loss occurs, whichever happens first). The above propositions slightly generalize the results by Besag and Clifford [3] about $Q^{\mathrm{BC}}_{\gamma(1,T)}$, because it can potentially stop at any arbitrary stopping time before $\gamma_{1,T}$.
Appendix D Nonstandard asymptotic behavior of log-optimal bets and wealth
In the following, we characterize the long-term behavior of the log-optimal strategy and the corresponding logarithmic wealth.
Theorem D.1.
Let $B_{t}^{*}$ be the log-optimal betting strategy (10) at step $t$. It holds that $B_{t}^{*}(I_{t})\to 1$ almost surely for $t\to\infty$.
This result means that even the log-optimal betting strategy will eventually have its wealth multiplied a factor arbitrarily close to one. This is a key aspect where our problem setting differs from many other standard i.i.d.-like problems — for the latter, the optimal strategies usually multiply the wealth of the gambler by a constant factor strictly larger than one in expectation. The reason for this unusual behavior is that in our case the indicators $I_{1},I_{2},\ldots$ are exchangeable under both, the null and the alternative. Therefore, the null and alternative hypothesis learn the same distribution of $I_{t}$ in the long run. Thus, even if $H_{1}$ was true, at some point we can longer exploit this knowledge by betting against $H_{0}$.
Proof.
From Section 2.2, we know that $I_{1},I_{2},\ldots$ are i.i.d. conditional on $Q_{\mathrm{lim}}$ with $\mathbb{E}(I_{1}|Q_{\mathrm{lim}})=Q_{\mathrm{lim}}$. Furthermore, for $p_{t}^{*}$ from Proposition 3.1, we have $p_{t}^{*}=\mathbb{E}(I_{t}|I_{1}^{t-1})=\mathbb{E}(I_{1}|I_{2}^{t})$, since $I_{1},\ldots,I_{t}$ are exchangeable. Levy’s zero-one law implies that $\mathbb{E}(I_{1}|I_{2}^{t})\stackrel{{\scriptstyle a.s.}}{{\to}}\mathbb{E}(I_{1}|I_{2}^{\infty})=\mathbb{E}(I_{1}|Q_{\mathrm{lim}},I_{2}^{\infty})=\mathbb{E}(I_{1}|Q_{\mathrm{lim}})=Q_{\mathrm{lim}}$ for $t\to\infty$ and thus $p_{t}^{*}\stackrel{{\scriptstyle a.s.}}{{\to}}Q_{\mathrm{lim}}$. We now show that $B_{t}^{*}(I_{t})|\{Q_{\mathrm{lim}}=q_{\mathrm{lim}}\}\stackrel{{\scriptstyle a.s.}}{{\to}}1$ for every possible realization $q_{\mathrm{lim}}$ of $Q_{\mathrm{lim}}$. First, consider the case $q_{\mathrm{lim}}\in(0,1)$. Then $p_{t}^{*}(t+1)/(L_{t-1}+1)|\{Q_{\mathrm{lim}}=q_{\mathrm{lim}}\}\stackrel{{\scriptstyle a.s.}}{{\to}}q_{\mathrm{lim}}\cdot 1/q_{\mathrm{lim}}=1$ and $(1-p_{t}^{*})(t+1)/(t-L_{t-1})|\{Q_{\mathrm{lim}}=q_{\mathrm{lim}}\}\stackrel{{\scriptstyle a.s.}}{{\to}}(1-q_{\mathrm{lim}})\cdot 1/(1-q_{\mathrm{lim}})=1$ for $t\to\infty$ and due to Proposition 3.1, we have $B_{t}^{*}(I_{t})|\{Q_{\mathrm{lim}}=q_{\mathrm{lim}}\}\stackrel{{\scriptstyle a.s.}}{{\to}}1$. Now consider the case $q_{\mathrm{lim}}=0$. Since $\mathbb{P}(I_{t}=1|Q_{\mathrm{lim}}=0)=0$ for all $t\in\mathbb{N}$, we have $L_{t}|\{Q_{\mathrm{lim}}=0\}=0$ for all $t\in\mathbb{N}$ almost surely. Since $(1-p_{t}^{*})(t+1)/(t-L_{t-1})|\{Q_{\mathrm{lim}}=0\}\stackrel{{\scriptstyle a.s.}}{{\to}}1$, we have $B_{t}^{*}(I_{t})|\{Q_{\mathrm{lim}}=0\}\stackrel{{\scriptstyle a.s.}}{{\to}}1$. If $q_{\mathrm{lim}}=1$, we analogously have $L_{t}|\{Q_{\mathrm{lim}}=1\}=t$ almost surely and $p_{t}^{*}(t+1)/(L_{t-1}+1)|\{Q_{\mathrm{lim}}=1\}\stackrel{{\scriptstyle a.s.}}{{\to}}1$. ∎
Corollary D.2.
Let $W_{T}^{*}$ be the wealth obtained by the log-optimal betting strategy after $T$ permutations. It holds $\log(W_{T}^{*})/T\to 0$ for $T\to\infty$ almost surely.
Again, for other i.i.d.-like problems in this area, the wealth grows exponentially and $\log(W_{T}^{*})/T$ usually tends to a constant larger than 0 under the alternative. In our problem however, the wealth (and its maximum) will hit a plateau, and at some point sampling more permutations does not help anymore (as would be expected).
Proof.
Let $\epsilon>0$. Due to Theorem D.1, there exists $N\in\mathbb{N}$ such that $|\log(B_{t}^{*}(I_{t}))|<\epsilon$ a.s. for all $t\geq N$. Therefore,
$|\log(W_{T}^{*})|\leq\sum_{t=1}^{N-1}|\log(B_{t}^{*}(I_{t}))|+\sum_{t=N}^{T}|\log(B_{t}^{*}(I_{t}))|<a+\epsilon T$
a.s. for all $T>N$, where $a$ is some non-negative constant. Hence, there exists $n\in\mathbb{N}$ such that $|\log(W_{T}^{*})/T|<\epsilon$ a.s. for all $T\geq n$. ∎
Theorem D.1 and Corollary D.2 show that at a late stage we can only gain wealth very slowly. Therefore, we need to bet using the structure of the alternative carefully in order to be able to build up a significant wealth in an appropriate time period. In the following section, we introduce a simple strategy based on these results.
Appendix E Using a Beta distribution as working prior for $Q_{\mathrm{lim}}$
@@
The function $\mathrm{Beta}(a,b)$ is called beta function. In this section we derive the wealth obtained by applying the binomial mixture strategy with a beta distribution as working prior for $Q_{\mathrm{lim}}$ and show some simulation results for this strategy.
Proposition E.1.
After $T$ permutations and $\ell$ losses, the wealth obtained by the binomial mixture strategy with density $g_{a,b}$ is given by $W_{T}^{g_{a,b}}(\ell)=\mathrm{Beta}(a+\ell,b+T-\ell)/(\mathrm{Beta}(\ell+1,T-\ell+1)\mathrm{Beta}(a,b))$.
@@
First note that $(T+1){T\choose\ell}=\mathrm{Beta}(\ell+1,T+1)$. Therefore, the wealth is given by
@@
$=\frac{\mathrm{Beta}(\ell+a,T+b-\ell)}{\mathrm{Beta}(\ell+1,T-\ell+1)\mathrm{Beta}(a,b)}.$
It can be shown that if $a\leq 1$ and $b\geq 1$, then $g_{a,b}(x)$ is nonincreasing in $x$ which seems reasonable as working prior for $Q_{\mathrm{lim}}$. In particular, for $a=1$ and $b\in\mathbb{N}$, the wealth simplifies to
$W_{T}^{g_{1,b}}(\ell)=b\frac{(T+1)\cdots(T-\ell+1)}{(T+b)\cdots(T-\ell+b)},$
which is decreasing for an increasing number of losses. In Figure 7 the log-transformed p-values obtained by the binomial mixture strategy with density $g_{1,b}$ are compared with the ones obtained by the binomial strategy, the binomial mixture strategy (with density $u_{0.9\alpha}$), the aggressive strategy and the classical permutation p-value. We used the same simulation setup as described in Section 7. The binomial mixture strategy with beta working prior bets more aggressively when $b$ is large. However, the binomial strategy and the binomial mixture strategy with uniform working prior lead to more rejections, as seen by the larger number of log-transformed p-values below the grey line $\log(\alpha)$, in all cases.
Figure 7: All plots display log-transformed p-values in ascending order for $2000$ simulations — following the experimental protocol in Subsection E. The grey horizontal line equals $\log(\alpha)$, where $\alpha$ equals 0.05. The binomial mixture strategy with a beta working prior bets more aggressive when $b$ is large. However, in terms of rejections (p-values below grey line) it is outperformed by the binomial strategy and the binomial mixture strategy with uniform working prior in all cases.
Appendix F Additional simulation results
In this section, we provide additional simulation results for the comparison of the Besag-Clifford method with our randomized binomial mixture strategy in terms of power and number of permutations. The simulation setup is the same as described in Section 7, however, in each figure we varied one of the parameters. In Figure 8, the observations follow a log-normal distribution (control observations were drawn from $\exp(Z)$ and treated observations from $\exp(Z+\mu)$, where $Z$ follows a standard normal distribution). In Figures 9 and 10, we chose the parameter of the binomial mixture strategy as $c=0.8\alpha$ and $c=0.99\alpha$, respectively. In Figures 1114, the sample size in each trial was varied ($100$, $200$, $500$ and $2000$).
All figures show the same behavior as in Section 7: the power of both methods is nearly identical, while the binomial mixture strategy reduces the number of permutations substantially.
Figure 8: Power and average number of permutations until the decision was obtained for log-normally distributed data — following the experimental protocol in Appendix F.
Figure 9: Power and average number of permutations until the decision was obtained for $c=0.8$ — following the experimental protocol in Appendix F.
Figure 10: Power and average number of permutations until the decision was obtained for $c=0.99$ — following the experimental protocol in Appendix F.
Figure 11: Power and average number of permutations until the decision was obtained for $n=100$ — following the experimental protocol in Appendix F.
Figure 12: Power and average number of permutations until the decision was obtained for $n=200$ — following the experimental protocol in Appendix F.
Figure 13: Power and average number of permutations until the decision was obtained for $n=500$ — following the experimental protocol in Appendix F.
Figure 14: Power and average number of permutations until the decision was obtained for $n=2000$ — following the experimental protocol in Appendix F.